# Reconstruction the scalar-torsion gravity version from the frame of exact cosmological solutions

### v3

@@ 4 Reconstruction of non-minimal coupling function from linear dependence $r\sim(1-n_{S})$
Thus, the gravitational term in action (1) for exact solutions (10)–(14) in quasi de Sitter regime $H\simeq const$ can be defined as
$f(T,\phi)_{STG}=-G(\phi)\sqrt{T}=cosnt\times\sqrt{T}\times\left[V(\phi)\right]^{-\frac{1+k}{2(m-1)}},$ (50)
i.e. the type of the non-minimal coupling of a scalar field $\phi$ with torsion scalar $T$ depends on the type of its potential.
After substitution of expression (49) into (47) we obtain
$H(t)=const\times\left[V(\phi)\right]^{-k/2}.$ (51)
Also, for $\alpha_{1}=-1$, from expression $\alpha_{1}=-\frac{r}{48\epsilon}$ we obtain the following condition on the value of the slow-roll parameter $\epsilon=r/48$ corresponding to the flat tensor spectrum, for $\epsilon>r/48$ one has red tilted tensor spectrum, and for $480<\epsilon<r/$ 48 one has blue tilted tensor spectrum. Thus, these inflationary models can predict different types of tensor perturbation spectrum depending on how close the cosmological dynamics of the early universe are to the purely exponential (de Sitter) expansion regime.
5 Inflation with exponential potential based on the scalar-torsion gravity $f(T,\phi)_{STG}=-G(\phi)\sqrt{T}$
As the example of cosmological model implying generalized scalar field potential (in the sense of positive and negative values), the exit from inflation and the second accelerated expansion of the universe we consider the model with exponential potential, which was considered earlier in the context of inflationary models based on GR, namely
$V(\phi)=V_{0}\exp(-2\beta\phi),$ (52)
where $\beta$ is the positive constant, and $V_{0}$ can take positive and negative values [54], that can be considered as consequence of the signature reversal symmetry [50].
From equations (11), (49) and (51) for potential (52) we obtain expressions for the Hubble parameter, coupling function, and kinetic function in explicit form
$H(t)=\lambda\exp\left(\beta k\phi\right),$ (53)
$G(\phi)=G_{0}\exp\left(\frac{\beta(k+1)}{m-1}\phi\right),$ (54)
$\omega(\phi)=\omega_{0}\exp\left(\frac{2\beta(m+k)}{m-1}\phi\right),$ (55)
where $\lambda$, $G_{0}$ and $\omega_{0}$ are a some constants.
Also, from expression (53) and following representation the Hubble parameter
$H(t)=\lambda+f(t),$ (56)
we obtain explicit relation between the type of the scalar field evolution and cosmological dynamics
$\phi(t)=\frac{1}{\beta k}\ln\left(1+\frac{f(t)}{\lambda}\right),$ (57)
where $f=f(t)$ is a some function of cosmic time.
Thus, in this case, we can construct exact solutions with arbitrary cosmological dynamics for first-order models based on the scalar-torsion gravity under consideration.
5.1 Exponential-power law inflation
As the example of the function corresponding to inflationary stage, the graceful exit from inflation and the second accelerated expansion of the universe at the large times we consider the case $f(t)=s/t$, where $s$ is a some constant.
The Hubble parameter and the scale factor in this case are
$H(t)=\lambda+\frac{s}{t},$ (58)
$a(t)\sim t^{s}\exp(\lambda t),$ (59)
where $s$ is a some positive constant, and $\dot{H}=-s/t^{2}<0$, i.e the Hubble parameter is the decreasing function of cosmic time.
Now, we define the e-fold number for this model as follows
$N(t)=\ln\left(\frac{a}{a_{0}}\right)=\lambda t+s\ln t,$ (60)
where $a_{0}$ is the scale factor at the beginning of inflation corresponding to $N=0$.
From expression (60) we can obtain the inverse dependence
$t(N)=\exp\left[\frac{N}{s}-W\left(\frac{\lambda}{s}e^{N/s}\right)\right],$ (61)
where $W$ is denote the Lambert function.
After substituting the dependence (61) into (57) and (52) we obtain the scalar field and the potential as the functions of the e-folds number
$\phi(N)=\frac{1}{\beta k}\ln\left\{1+\frac{s}{\lambda}\exp\left[W\left(\frac{\lambda}{s}e^{N/s}\right)-\frac{N}{s}\right]\right\},$ (62)
$V(N)=V_{0}\left\{1+\frac{s}{\lambda}\exp\left[W\left(\frac{\lambda}{s}e^{N/s}\right)-\frac{N}{s}\right]\right\}^{-2/k}.$ (63)
As one can see, the scalar field and potential at the beginning of inflation $\phi_{i}=\phi(N=0)$, $V_{i}=V(N=0)$ and at the end of inflation $\phi_{e}=\phi(N\approx 60)$, $V_{e}=V(N\approx 60)$ can take on different values due to the fact that they are determined by the five constant parameters of this model $\beta$, $k$, $s$, $\lambda$ and $V_{0}$.
We also note that the energy scale of inflation, which is defined as $E_{f}=|V|^{1/4}$ [69, 70] can be expressed as $E_{f}=|V_{0}e^{-2\beta\phi}|^{1/4}\sim 10^{-2}$ in chosen system of units or $E_{f}=|V_{0}e^{-2\beta\phi}|^{1/4}\sim 10^{-2}M_{P}\sim 10^{16}$ GeV [1, 69, 70] gives the additional relation between constant model’s parameters.
A detailed analysis of cosmological dynamics, within the framework of which it was shown that scale factor (59) implies an accelerated inflationary expansion of the universe, exit from inflation and second inflation (with a much lower rate of accelerated expansion of the universe than during the first inflation) at large times, was presented in [58, 59, 60, 61, 62].
We only note the general property of models with function $f=f(t)$ decreasing in time (as for the case of the function $f(t)=s/t$).
If condition $f(t\rightarrow\infty)=0$ is met, then from expressions (52)–(57) at the large times we have $\phi=0$, $H=\lambda$, $G=G_{0}$, $\omega=\omega_{0}$ and $V=V_{0}$.
Thus, at large times one has $f(t\gg t_{inf})\ll\lambda$ and $H\simeq\lambda$, and from (20) we obtain the model closed to the teleparallel equivalent of general relativity
$f(T,\phi)\simeq-\frac{G_{0}}{2\sqrt{T_{0}}}T+\alpha_{2}\left(V_{0}+\frac{1}{2}G_{0}\sqrt{T_{0}}\right)=$
$-\frac{1}{2}T+\alpha_{2}\left(V_{0}+\frac{1}{2}G^{2}_{0}\right)=-\frac{1}{2}T-\Lambda,$ (64)
with the cosmological constant $\Lambda=-\alpha_{2}\left(V_{0}+\frac{1}{2}G^{2}_{0}\right)$ for $G_{0}=\sqrt{T_{0}}$.
We also note that taking into account the observational value of the cosmological constant, we can consider the negative potential $V_{0}<0$ (and $\alpha_{2}<0$ due to condition (34), for the partial case $\alpha_{2}=-1$) to obtain $\Lambda=-\alpha_{2}\left(V_{0}+\frac{1}{2}G^{2}_{0}\right)=(2.846\pm 0.076)\times 10^{-122}$ [27].
Obviously, to build a correct model of the post-inflationary universe, it is necessary to take into account additional material fields, in particular, baryonic matter, radiation and Dark Matter. However, since we have a cosmological constant at the large times only, for the Cold Dark Matter this model at present era can be reduces to one close to the $\Lambda$CDM model based on TEGR.
5.2 Cyclic universe
We also note that based on the proposed approach, it is possible to obtain exact solutions for an arbitrary types of cosmological dynamics, for example, for a cyclic universe or Quasi-Steady State cosmological model, with the following Hubble parameter [63, 64, 65, 66]
$H(t)=\lambda-\frac{\mu\alpha\sin(\mu t)}{1+\alpha\cos(\mu t)},~{}~{}~{}f(t)=-\frac{\mu\alpha\sin(\mu t)}{1+\alpha\cos(\mu t)},$ (65)
and scale factor
$a(t)\sim\exp(\lambda t)\left(1+\alpha\cos(\mu t)\right),$ (66)
where constant $\mu>0$ and $10\leq\alpha<$ 1 in order to obtain the positive scale factor $a(t)>0$ for any cosmic time.
In this case, the Hubble parameter might increase at late times to account the well-known tension between PLANCK and local observations [55, 56, 57].
Corresponding scalar field evolution (57) for the Hubble parameter (65) is
$\phi(t)=\frac{1}{\beta k}\ln\left[1-\frac{\mu\alpha\sin(\mu t)}{\lambda(1+\alpha\cos(\mu t))}\right].$ (67)
In this case, the potential (52), coupling function (54) and kinetic function (55) can be written as follows
$V(t)=V_{0}\left[1-\frac{\mu\alpha\sin(\mu t)}{\lambda(1+\alpha\cos(\mu t))}\right]^{-2/k},$ (68)
$G(t)=G_{0}\left[1-\frac{\mu\alpha\sin(\mu t)}{\lambda(1+\alpha\cos(\mu t))}\right]^{\frac{k+1}{k(m-1)}},$ (69)
$\omega(t)=\omega_{0}\left[1-\frac{\mu\alpha\sin(\mu t)}{\lambda(1+\alpha\cos(\mu t))}\right]^{\frac{2(m+k)}{k(m-1)}},$ (70)
i.e. they are the periodic functions of cosmic time, and model is reduced to the case of cosmological constant $\Lambda$ and TEGR at the times closed to $t_{\Lambda}=\pi n/\mu$, where $n$ are a some positive integer.
Thus, the Hubble tension in this case can be explained by the deviation of the observation time in present era from this value $t_{obs}=\frac{\pi}{\mu}(n+\Delta)$.
The solution to the Hubble tension problem based on the Quasi-Steady State cosmological model based on background dynamics without analyzing the model’s compliance with observational constraints on the anisotropy and polarization of CMB was discussed earlier in [65, 66]. In the case of the proposed model, these constraints are satisfied due to relation (45). Other possible solutions to the Hubble tension problem can be found, for example, in [55, 56, 57].
Obviously, the cosmological models with other scalar field potentials and different types of cosmological dynamics will also satisfy observational constraints on the values of the parameters of cosmological perturbations due to dependence (45) for the proposed gravity theory.
6 The properties of the proposed models
As one can see, different cosmological dynamics does not affect the linear relation $r\sim(1-n_{S})$ between tensor-to scalar ratio and spectral tilt of the scalar perturbations (45), i.e. the cosmological models under consideration with different dynamics of accelerated expansion of the early universe can be considered as verifiable ones by the observational constraints on the parameters of cosmological perturbations.
Also, an obvious generalization of the proposed approach is the analysis of cosmological models based on the other scalar field potentials which differ from exponential one (52) based on expressions (11), (49) and (51) or for representations of the coupling function other than (49).
Thus, the proposed scalar-torsion gravity $f(T,\phi)_{STG}\sim G(\phi)\sqrt{T}$ allows one to consider inflationary models with an arbitrary potential of the scalar field and arbitrary cosmological dynamics (satisfying the correct change of stages of expansion of the universe) as in principle verifiable by observational constraints already at the first (linear) order of expansion of dependence $r=r(1-n_{S})$.
Let us note the results presented in [34], in which the verification of inflationary models based on scalar-tensor gravity also did not depend on the type of scalar field potential or the specific type of scalar-tensor gravity, however, in this case, the quadratic dependence $r\sim(1-n_{S})^{2}$ corresponding to the second-order models within the framework of our classification was obtained.
However, we note that verification of inflationary models with different dependencies $r=r(1-n_{S})$ with specific model’s parameters can be achieved as for the case of GR [48] and for the case of the other modified gravity theories as well (see, for example, in [29, 30, 31, 32, 33]).
For these reasons, the proposed approach is one of the many possible ways to construct the cosmological models. Nevertheless, the main advantage of the proposed approach is the ability to obtain the verified cosmological models for an arbitrary model’s parameters (including arbitrary dynamics of the expansion of the universe) for the exactly defined type of the scalar-torsion gravity. This property of the proposed theory of gravity that makes it attractive for further research and applications to actual problems of cosmology and astrophysics.
7 Conclusion
In this work, cosmological models based on scalar-torsion gravity with non-minimal coupling between the scalar field and torsion, which was obtained as consequence of the generalized exact solutions of cosmological dynamic equations for an arbitrary Hubble parameter $H=H(t)$ and an arbitrary evolution of a scalar field $\phi=\phi(t)$.
Based on these generalized exact solutions, it was concluded that the gravitational interaction within the framework of scalar-torsion gravity in such models should be determined by a function $f(T,\phi)_{STG}\sim G(\phi)\sqrt{T}$, which corresponds to some corrections to the usual teleparallel equivalent of general relativity $f(T,\phi)_{TEGR}\sim T$. It was also noted that at the inflationary quasi de Sitter stage of the evolution of the universe, these corrections are quite small.
Further, based on the previously proposed approach to analyzing the evolution of cosmological perturbations for the cosmological models based on the scalar-torsion gravity for general case $f(T,\phi)_{STG}=f(T,\phi)$, we determined the parameters of cosmological perturbations for the case of the considered version of scalar-tensor gravity $f(T,\phi)_{STG}\sim G(\phi)\sqrt{T}$ on the basis of exact background solutions of cosmological dynamic equations.
A classification of inflation models according to the order of expansion of the dependence tensor-to-scalar ratio from spectral index of the scalar perturbations $r=r(1-n_{S})$ was also proposed, and on the basis of this classification the method for constructing verified by observational constraints inflationary models based on scalar-torsion gravity $f(T,\phi)_{STG}=-G(\phi)\sqrt{T}$ was considered, implying a linear dependence $r\sim(1-n_{S})$ for arbitrary model’s parameters with specific relation (46) between the Hubble parameter and the coupling function. We note that in this case, the type of scalar field potential $V=V(\phi)$ or other background parameters do not affect the possibility of verifying the proposed cosmological models.
It was also noted that proposed inflationary models can predict different types of tensor perturbation spectrum (red, blue of flat) depending on the value o the tensor-to-scalar ratio and how close the cosmological dynamics of the early universe are to the purely exponential (de Sitter) expansion regime.
Also, due to the correspondence of arbitrary models of cosmological inflation based on the generalized exact solutions to observational constraints on the parameters of cosmological perturbations, the following relationship between the non-minimal coupling function and the potential of the scalar field $G(\phi)\sim\left[V(\phi)\right]^{-\frac{1+k}{2(m-1)}}$ for quasi de Sitter dynamics accelerated expansion of the universe with the Hubble parameter $H\simeq const$ was obtained. Thus, the type of non-minimal coupling between the scalar field and torsion depends on the type of the scalar field potential.
As an example, we consider cosmological models based on exponential potential. It was shown that by choosing the Hubble parameter $H=H(t)$ it is possible to construct models with an exit from inflation and the second accelerated expansion of the universe in the present era. Also, as a possible solution to the Hubble tension problem, a model of a cyclic universe was considered on the basis of the proposed scalar-torsion gravity theory.
In conclusion, we note that proposed type of the scalar-torsion gravity $f(T,\phi)_{STG}\sim G(\phi)\sqrt{T}$ reconstructed from exact solutions of cosmological dynamic equations and implying the wide class of verified cosmological models with arbitrary parameters is of interest for the further research of the deviations in the spectrum of relict gravitational waves and in the evolution of compact astrophysical objects from teleparallel equivalent of general relativity or from the other modified gravity theories.

### v4

@@ IV Reconstruction of the parameters of verified inflationary models
Therefore, the gravitational term in action (1) for exact relations (10)–(12) in quasi de Sitter regime $H\simeq const$ can be defined as
$f(T,\phi)_{STG}=-G(\phi)\sqrt{T}\sim\sqrt{T}\times\left[V(\phi)\right]^{-\frac{1+k}{2(m-1)}}.$ (44)
Thus, for any physical potential of a scalar field Martin:2013tda , from expression (44) one can reconstruct the type of scalar-torsion gravity corresponding to the possibility of verifying different realizations of the inflationary scenario.
V Conclusion
In this work, cosmological models based on the certain type of the scalar-torsion gravity $f(T,\phi)_{STG}\sim G(\phi)\sqrt{T}$ are considered. This type of the scalar-torsion gravity was obtained as consequence of the generalized exact solutions of cosmological dynamic equations.
A classification of inflation models according to the order of expansion of the dependence tensor-to-scalar ratio from spectral index of the scalar perturbations $r=r(1-n_{S})$ was also proposed. On the basis of this classification the method for constructing verified inflationary models based on scalar-torsion gravity $f(T,\phi)_{STG}=-G(\phi)\sqrt{T}$ for the first-order models $r\sim(1-n_{S})$ was considered. In this case, the type of scalar field potential $V=V(\phi)$ or other background parameters don’t affect the possibility of verifying the proposed cosmological models.
Also, due to the correspondence of arbitrary models of cosmological inflation based on the generalized exact solutions to observational constraints on the parameters of cosmological perturbations, the following relationship between the non-minimal coupling function and the potential of the scalar field $G(\phi)\sim\left[V(\phi)\right]^{-\frac{1+k}{2(m-1)}}$ for quasi de Sitter dynamics accelerated expansion of the universe was obtained. Thus, the type of non-minimal coupling between the scalar field and torsion depends on the type of the scalar field potential for the proposed inflationary models.
In conclusion, we note that proposed type of the scalar-torsion gravity $f(T,\phi)_{STG}\sim G(\phi)\sqrt{T}$ reconstructed from exact solutions of cosmological dynamic equations and corresponding wide class of verified cosmological models with arbitrary parameters is of interest for the further research of the deviations in the spectrum of relict gravitational waves from teleparallel equivalent of general relativity or from the other modified gravity theories.